NOAA SATOPS Morning Report: Friday, January 04, 2013


SATOPS Morning Report: Friday, January 04, 2013

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 004

GOES
All GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
OPS54 DOY004 07:15z data loss of 5811 s/c seconds due to a ground system problem at PEAK.

OPS51 DOY004 08:43z rev 31822 lost due to a LION ground station problem.

All other DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
F-16 DMSP/ SSMI (Special Sensor Microwave/Image) Data, rev. 47542 was lost per FNMOC (Fleet Numerical Meteorological and Oceanographic Center).

F-17 DMSP/ SSMI (Special Sensor Microwave/Image) Data, rev. 31822 was lost per FNMOC (Fleet Numerical Meteorological and Oceanographic Center).

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.